OpenStudy (anonymous):

Find the nth term of the arithmetic sequence where a1 = 5000 and d = -100.

jimthompson5910 (jim_thompson5910):

The nth term of any arithmetic sequence is an = a1 + (n-1)*d In this case a1 = 5000 and d = -100, so an = a1 + (n-1)*d an = 5000 + (n-1)*(-100) Simplify this to get your answer

OpenStudy (anonymous):

= -100n + 5100.

jimthompson5910 (jim_thompson5910):

you got it